Determine whether or not the following matrices are in row echelon form or not? (4.1)

(1 2 −2)

( 0 1 2)

( 0 0 1)


(4.2) (1 2 −2)

(0 1 2 )

( 0 0 1 ) 


1
Expert's answer
2021-05-25T17:25:13-0400

A matrix is in row echelon form (ref) when it satisfies the following conditions.

The first non-zero element in each row, called the leading entry, is 1.

Each leading entry is in a column to the right of the leading entry in the previous row.

Rows with all zero elements, if any, are below rows having a non-zero element.


4.1 Yes. The matrix


"\begin{pmatrix}
 1 & 2 &-2 \\
 0 & 1 & 2 \\
0 & 0 & 1
\end{pmatrix}"


is in row echelon form, since matrix satisfies all of the requirements for a row echelon matrix.


But the matrix is not in reduced row echelon matrix.

"R_1=R_1-2R_2"


"\begin{pmatrix}
 1 & 0 & -6 \\
 0 & 1 & 2 \\
0 & 0 & 1
\end{pmatrix}"

"R_1=R_1+6R_3"


"\begin{pmatrix}
 1 & 0 & 0 \\
 0 & 1 & 2 \\
0 & 0 & 1
\end{pmatrix}"

"R_2=R_2-2R_3"

"\begin{pmatrix}
 1 & 0 & 0 \\
 0 & 1 & 0 \\
0 & 0 & 1
\end{pmatrix}"





4.2 Yes. The matrix


"\begin{pmatrix}
 1 & 2 &-2 \\
 0 & 1 & 2 \\
0 & 0 & 1
\end{pmatrix}"


is in row echelon form, since matrix satisfies all of the requirements for a row echelon matrix.


But the matrix is not in reduced row echelon matrix.

"R_1=R_1-2R_2"

"\begin{pmatrix}
 1 & 0 & -6 \\
 0 & 1 & 2 \\
0 & 0 & 1
\end{pmatrix}"

"R_1=R_1+6R_3"

"\begin{pmatrix}
 1 & 0 & 0 \\
 0 & 1 & 2 \\
0 & 0 & 1
\end{pmatrix}"

"R_2=R_2-2R_3"

"\begin{pmatrix}
 1 & 0 & 0 \\
 0 & 1 & 0 \\
0 & 0 & 1
\end{pmatrix}"